# Colloquium: Synthetic quantum matter in non-standard geometries

### v1

@@ III Applications - novel phenomena in exotic geometries
This section of the colloquium shall provide an overview of the fascinating topics that can be addressed by implementing the different types of exotic lattices discussed in the previous section. These phenomena interest researchers from various fields, ranging from condensed matter physics to cosmology.
@@ III.1.1 Single-particle localization
Anderson [11] demonstrated that single electrons in uncorrelated static disordered materials exhibit vanishing diffusion at long times, with an exponentially localized wavefunction. The key mechanism in 1D and 2D systems is destructive wave interference everywhere except at the initial point, leading to inevitable localization even in the presence of infinitesimal disorder. However, in 3D, complete backscattering is less likely, requiring a stronger disorder strength for localization to occur, and giving rise to a metal-insulator transition at a critical disorder strength. The transition is characterized by a power-law divergence of the localization length near the critical energy, that is, the mobility edge. In recent times, the phenomenon of localization has also been pursued in the context of quasiperiodic potentials, revealing a rich array of phenomena that challenge our conventional understanding. As compared to disordered systems, quasiperiodic systems display a more diverse localization behavior, including multifractal states and mobility edges even in one dimension. This diversity arises from the unique interplay between disorder and the inherent periodicity of quasiperiodic structures, providing an alternative perspective on the fundamental mechanisms of localization.
@@
with being the amplitude of the state at site , and being the total number of sites in the system. The mean inverse participation ratio is obtained by averaging over all eigenstates, . For a perfectly localized system, with all states pinned to a single specific site, becomes 1, whereas for perfectly extended (delocalized) states . Hence, the scaling of with is a practically useful indicator to differentiate localized/extended states. Another classification tool is the -fractal dimension [97, 78] of the eigenstates. For a specific state, this quantity is defined by the following relation . For , the -fractal dimension is proportional to the logarithm of the inverse participation ratio of a state. Averaging over all eigenstates, the mean fractal dimension approaches the value 1 (0) for extended (localized) systems, whereas intermediate values suggest multifractal behavior. The information carried by the -fractal dimensions can also be expressed through the multifractal spectrum , where and [97].
@@
Following Boers et al. [35], Li et al. [144], a 1D quasicrystal with an energy-dependent mobility edge has been implemented with cold atoms in an optical lattice, see Lüschen et al. [154]. An alternative atomic realization of the generalized AA model has been presented in An et al. [7] using a momentum-space lattice.
@@
In contrast to the AA model, the Fibonacci chain (see Sec. I) does not exhibit a localization transition and remains critical [180, 126] for any value of the modulation strength as dictated by the Luck criterion of bounded fluctuations in the pattern [151]. While the fluctuations can become arbitrarily large for systems with uncorrelated noise, for many quasicrystals, including the Fibonacci quasicrystal, fluctuations are bounded. Systems with bounded and unbounded fluctuations have been shown to belong to different universality classes. Using transfer matrix method [24], scaling analysis [151], real-space renormalization group [110, 4], the critical behavior of quantum Ising chains with Fibonacci modulation has been studied, establishing conditions for the occurrence of magnetic phase transitions, and for the quasiperiodic system to be in a different universality class than the uniform one [151, 4]. Also the critical behavior of the Heisenberg model, with quasiperiodic couplings has been studied [103, 104, 111, 4], as well as the XXZ model [256]. It has been shown that a Fibonacci modulation of the Heisenberg coupling alters its critical behavior, in contrast to the XY model or the quantum Ising model, where a quasiperiodic modulation of the coupling is marginal or irrelevant as the central charge remains unaffected. However, we come to these later since these spin models map to interacting fermionic models.
@@
Non-interacting systems in lower-dimensional landscapes, defined by static, random potential energy and short-ranged tunneling, exhibit Anderson localization. Transitioning to higher dimensions facilitates delocalization, driven by increased energy density. For quasicrystals, with the increasing sophistication of experimental techniques, research into localization has extended to two-dimensional (2D) quasicrystals, revealing a multifaceted landscape of localization behavior. Quasicrystalline order usually arises from projecting a higher-dimensional periodic lattice in an incommensurate manner. Penrose introduced a method for constructing quasicrystals using a set of tiles and matching rules, resulting in the well-known fivefold symmetric Penrose tiling and the eightfold symmetric octagonal tiling [186]. These tilings, with disallowed rotational symmetries, exhibit self-similarity in both real and reciprocal space. Self-similarity upon scaling indicates nontrivial structure at arbitrarily large scales, reflected in quasicrystal diffraction patterns with sharp peaks at very small momenta. In the studies Viebahn et al. [255], Sbroscia et al. [212], a BEC of Potassium atoms is utilized to probe a 2D quasicrystalline optical lattice through stimulated two-photon Kapitza-Dirac scattering. The quasicrystalline potential is created by combining four mutually incoherent 1D optical lattices, forming a global eightfold rotational symmetry. Unlike periodic lattices, the combination of reciprocal lattice vectors results in new, smaller momentum scales, creating a self-similar fractal structure (see Fig. 1b). This process is akin to the incommensurate projection of a 4D simple cubic lattice to the 2D plane, similar to the construction of the octagonal tiling. Adjusting the number of lattice beams allows control over the dimensionality of the parent lattice. Finally, adiabatically loading BECs into quasicrystal lattices allows studying Bose glasses in these geometries [285].
The system thermalizes at higher energy density and weaker disorder, exhibiting a conducting phase.
@@ III.1.2 Many-body localization
Quantum simulators of quasicrystals have also yielded groundbreaking discoveries in the many-body context, showcasing their unparalleled ability to test theoretical predictions experimentally. Traditionally, a disordered potential localizes single-particle quantum eigenstates, rendering the system an insulator with zero conductivity. In Basko et al. [21], it was proposed that interacting many-body systems could similarly undergo a MBL transition. This transition involves highly excited states, occurring at low energy density and strong disorder, where the system remains a perfect insulator. The system thermalizes at higher energy density and weaker disorder, exhibiting a conducting phase. MBL violates the eigenstate thermalization hypothesis (ETH), which posits that each eigenstate at finite energy density has the same expectation values of local physical observables as the ones provided by local thermal states. This implies that thermalization occurs at the level of individual eigenstates in an isolated quantum system, where no memory of the initial state survives. The eigenstates of the system in the MBL phase do not obey the ETH, and some memory of the initial conditions is retained in local observables for arbitrarily long times.
@@
In general, the theoretical description of the MBL transition is extremely challenging. In contrast to the metal-insulator transitions, which occur in the ground state or the low-energy states of a system, the MBL transition affects the whole spectrum or at least an extended spectral range, which might be far away from the ground state. For this reason, much of the machinery to study many-body phases of matter (such as Monte Carlo techniques or tensor network methods) is not well suited for studying MBL. As a consequence, numerical explorations have mostly relied on exact diagonalization studies of systems with 20-30 spins, and even these relatively moderately sized studies strongly profit from sophisticated techniques [188, 224], exploiting the sparseness of the matrix. Larger system sizes have been treated with a self-consistent Hartree-Fock description [270] or time-dependent density matrix renormalization group (DMRG) studies [140, 278].
Different fingerprints of the MBL phase are known: (i) The half-chain von Neumann entanglement entropy is strongly suppressed in the MBL phase [138, 140, 156]. The transition point can be estimated by considering the value at which the entanglement entropy per site is independent of the system size . (ii) The level spacing ratio , where is the spacing between adjacent energy levels, differentiates between localized and ergodic phase. It is well established by random matrix theory that the level spacing distribution characterizes the statistical ensemble to which a matrix belongs [96]. Ergodic systems typically belong to the Wigner-Dyson ensemble, characterized by level repulsion and a relatively large value . On the other hand, integrable/localized systems are characterized by the Poisson statistics, , see Pal and Huse [184]. (iii) Participation ratio and fractal dimension of the eigenstates, as introduced in the Section on non-interacting systems, can also be used for a many-body system if, instead of the full many-body eigenstates, the eigenstates of the one-body density matrix are considered [25, 156]. (iv) Dynamics probes have been used in many experiments: If a system is localized, it should keep memory of its initial state during a quench. Often, systems are prepared in a state with alternating empty and occupied sites, and the evolution of this even/odd population imbalance is tracked. In contrast to an ergodic system, the MBL system remains imbalanced on any time scale [216].
@@
Given the computational difficulties, quantum simulation becomes an extremely valuable tool to study MBL. At the same time, due to the instability of the MBL phase against a thermal bath, the study of the MBL phase is also motivated by the unique opportunity of isolating quantum systems from the environment offered by quantum simulators. The primary systems where the MBL transition is experimentally studied are cold atoms in optical lattices with an AA modulation [216, 153, 125, 152, 199]. Experiments in Munich [216, 153, 125] have concentrated on rather larger systems, arrays of 1D tubes with approximately 150-200 sites per tube, filled with spin-1/2 fermions.
@@
An important figure of merit in these experiments is the population imbalance between even and odd sites. Initially, the system was prepared for a large imbalance, which was then washed out through dynamic evolution. While in the absence of a quasiperiodic potential, the imbalance fully disappears on short time scales, a sufficiently strong quasiperiodic potential leads to saturation of the imbalance at a finite value, see Fig. 9(a). By plotting the saturation value versus the potential strength , see Fig. 9(b), the onset of localization is found to be shifted towards larger values of by the interactions. In the non-interacting case, the transition seems to occur slightly below the single-body critical value , which can be accounted for by the harmonic trapping potential by comparison with exact numerical simulations. In the vicinity of the localization transition, especially in the interacting system, the saturation occurs only on much longer time scales than deeply in the ergodic or the MBL regime. Kohlert et al. [125] also studied MBL in a GAA model. In contrast to the AA model, the non-interacting GAA model exhibits an intermediate phase in which both localized and delocalized eigenstates co-exist. While the experiment confirmed the MBL behavior of the interacting GAA, a many-body counterpart of the intermediate phase has not been seen.
@@
The experiments in Harvard, in particular Rispoli et al. [199], have concentrated on this critical regime, measuring observables like correlations between sites, entropy, and particle number fluctuations. The experimental procedure of Lukin et al. [152], Rispoli et al. [199] starts from a bosonic Mott insulator with one atom per site and measures the particle number per site after some evolution time. From this, different quantities can be computed. For instance, the probability of finding particles on a site determines the single-site von Neumann entropy, . For weak quasiperiodic potentials, the obtained value matches well with the expected thermal average, but it becomes reduced when the quasiperiodic modulation becomes stronger, approximately for , see Fig. 9(c), notably beyond the critical value of the single-particle AA model. In contrast, the entropy reaches a maximum within the transition regime as the system strongly fluctuates. Criticality is also characterized by strong -site correlations , shown in Fig. 9(d).
@@
A bosonic system has also been produced with 9 superconducting qubits, realizing a 9-site Bose-Hubbard chain with an AA-like potential [203]. If prepared with only one or two bosons, all energy levels can be detected in this setup. In addition to obtaining the fractal butterfly spectrum in the single-particle case, this setup has also demonstrated that the level spacing of an interacting system is shifted towards smaller values by increasing the quasiperiodic modulation strength. Quantum simulators can also address bosonic systems and, therefore, help to solve the question of whether MBL exists in bosonic systems [55].
@@ III.2 Topological phenomena
Topological quantum systems are characterized by a band structure with non-trivial topological properties, as in the case of topological insulators [195], or by a topological many-body ground state, as in the case of FQH systems and topological superconductors [171]. In all cases, the non-trivial topological properties are mathematically defined via non-local quantities, specifically non-zero Chern numbers, which are physically linked to robust properties, such as transport behavior or fractional quantum statistics of quasiparticles. Topological matter has been an intense research field in the past decades, mostly in the context of systems in “regular” geometries and driven by the search for robust quantum systems. Quasicrystalline or fractal structures and higher-dimensional lattices provide interesting new aspects to the field of topological matter, as discussed in the following.

### v2

@@ III Applications - novel phenomena in exotic geometries
This colloquium section shall provide an overview of the fascinating topics that can be addressed by implementing the different types of exotic lattices discussed in the previous section. These phenomena interest researchers from various fields, from condensed matter physics to cosmology.
@@ III.1.1 Single-particle localization
Anderson [12] demonstrated that single electrons in uncorrelated static disordered materials exhibit vanishing diffusion at long times, with an exponentially localized wavefunction. The key mechanism in 1D and 2D systems is destructive wave interference everywhere except at the initial point, leading to inevitable localization even in the presence of infinitesimal disorder. However, in 3D, complete backscattering is less likely, requiring a stronger disorder strength for localization to occur, and giving rise to a metal-insulator transition at a critical disorder strength. The transition is characterized by a power-law divergence of the localization length near the critical energy, that is, the mobility edge. In recent times, the phenomenon of localization has also been pursued in the context of quasiperiodic potentials, revealing a rich array of phenomena that challenge our conventional understanding. Compared to disordered systems, quasiperiodic systems display a more diverse localization behavior, including multifractal states and mobility edges even in one dimension. This diversity arises from the unique interplay between disorder and the inherent periodicity of quasiperiodic structures, providing an alternative perspective on the fundamental mechanisms of localization.
@@
with being the amplitude of the state at site , and being the total number of sites in the system. The mean inverse participation ratio is obtained by averaging over all eigenstates, . For a perfectly localized system, with all states pinned to a single specific site, becomes 1, whereas for perfectly extended (delocalized) states . Hence, the scaling of with is a useful indicator for differentiating localized/extended states. Another classification tool is the -fractal dimension [106, 84] of the eigenstates. For a specific state, this quantity is defined by the following relation . For , the -fractal dimension is proportional to the logarithm of the inverse participation ratio of a state. Averaging over all eigenstates, the mean fractal dimension approaches the value 1 (0) for extended (localized) systems, whereas intermediate values suggest multifractal behavior. The information carried by the -fractal dimensions can also be expressed through the multifractal spectrum , where and [106].
@@
Following Boers et al. [36], Li et al. [158], a 1D quasicrystal with an energy-dependent mobility edge has been implemented with cold atoms in an optical lattice, see Lüschen et al. [170]. An alternative atomic realization of the generalized AA model has been presented in An et al. [8] using a momentum-space lattice. It has also been shown that a balanced bichromatic optical lattice, where the amplitude of both frequency components are comparable, leads to a model with a mobility edge, beyond the standard AA model.
@@
In contrast to the AA model, the Fibonacci chain (see Sec. I) does not exhibit a localization transition and remains critical [196, 138] for any value of the modulation strength as dictated by the Luck criterion of bounded fluctuations in the pattern [166]. While the fluctuations can become arbitrarily large for systems with uncorrelated noise, fluctuations are bounded for many quasicrystals, including the Fibonacci quasicrystal. Systems with bounded and unbounded fluctuations have been shown to belong to different universality classes. Using transfer matrix method [25], scaling analysis [166], real-space renormalization group [120, 4], the critical behavior of quantum Ising chains with Fibonacci modulation has been studied, establishing conditions for the occurrence of magnetic phase transitions, and for the quasiperiodic system to be in a different universality class than the uniform one [166, 4]. Also the critical behavior of the Heisenberg model, with quasiperiodic couplings has been studied [113, 114, 121, 4], as well as the XXZ model [277]. It has been shown that a Fibonacci modulation of the Heisenberg coupling alters its critical behavior, in contrast to the XY model or the quantum Ising model, where a quasiperiodic modulation of the coupling is marginal or irrelevant as the central charge remains unaffected. However, we come to these later since these spin models map to interacting fermionic models.
@@
Non-interacting systems in lower-dimensional landscapes, defined by static, random potential energy and short-ranged tunneling, exhibit Anderson localization. Transitioning to higher dimensions facilitates delocalization, driven by increased energy density. For quasicrystals, with the increasing sophistication of experimental techniques, research into localization has extended to two-dimensional (2D) quasicrystals, revealing a multifaceted landscape of localization behavior. Quasicrystalline order usually arises from projecting a higher-dimensional periodic lattice in an incommensurate manner. Penrose introduced a method for constructing quasicrystals using a set of tiles and matching rules, resulting in the well-known fivefold symmetric Penrose tiling and the eightfold symmetric octagonal tiling [202]. These tilings, with disallowed rotational symmetries, exhibit self-similarity in both real and reciprocal space. Self-similarity upon scaling indicates nontrivial structure at arbitrarily large scales, reflected in quasicrystal diffraction patterns with sharp peaks at very small momenta. In the studies Viebahn et al. [276], Sbroscia et al. [228], a BEC of Potassium atoms is utilized to probe a 2D quasicrystalline optical lattice through stimulated two-photon Kapitza-Dirac scattering. The quasicrystalline potential is created by combining four mutually incoherent 1D optical lattices, forming a global eightfold rotational symmetry. Unlike periodic lattices, the combination of reciprocal lattice vectors results in new, smaller momentum scales, creating a self-similar fractal structure (see Fig. 1b). This process is akin to the incommensurate projection of a 4D simple cubic lattice to the 2D plane, similar to the construction of the octagonal tiling. Adjusting the number of lattice beams allows control over the dimensionality of the parent lattice. It has been predicted that the spectrum consists of a localized low-energy part, an extended high-energy part, and an intermediate regime, in which localized and critical states alternate [318]. Finally, adiabatically loading BECs into quasicrystal lattices allows studying Bose glasses in this geometries [308]. The system thermalizes at higher energy density and weaker disorder, exhibiting a conducting phase.
Localization in fractal lattices.
Also, fractal structures, such as the Sierpiński gasket, are known to host localized eigenstates [73], in addition to an infinite number of extended states [286]. This can give rise to complex transport behavior, featuring an increased return probability in quantum walks as compared to classical random walks [63]. Recently, quantum transport in fractal geometries has been explored in waveguide experiments [301], and super-diffusive quantum transport has been observed in both Sierpiński gasket and carpet. However, in the experimental setup connectivity between all nearest neighbors has avoided a bottleneck in the Sierpiński gasket, which occurs at sites where two fractal generations are glued together. As pointed out by Ref. Rojo-Francàs et al. [218], this bottleneck significantly affects the energy spectrum of the system, and slows down quantum transport into the sub-diffusive regime. Theoretical studies of fractal lattices with random connectivity have shown that eigenfunction localization depends on the spectral dimension rather than the Hausdorff dimension [143].
Localization in higher-dimensional or curved spaces.
While quasicrystalline or fractal geometries may give rise to localization in the absence of disorder, quantum particles in Euclidean 1D or 2D lattices are localized by an infinitesimal amount of disorder. A different behavior has been reported for hyperbolic lattices, where finite disorder is required for localization [52], similar to the localization behavior in higher dimensions , see [261]. Quantum funneling, a localization effect caused by the singularity in a negatively curved space, has been described by Zhang et al. [312]. In the future, quantum simulators of exotic geometries may be used to probe localization behavior in higher-dimensional or curved spaces.
@@ III.1.2 Many-body localization
Quantum simulators of quasicrystals have also yielded groundbreaking discoveries in the many-body context, showcasing their ability to test theoretical predictions experimentally. Traditionally, a disordered potential localizes single-particle quantum eigenstates, rendering the system an insulator with zero conductivity. In Basko et al. [22], it was proposed that interacting many-body systems could similarly undergo a MBL transition. This transition involves highly excited states, occurring at low energy density and strong disorder, where the system remains a perfect insulator. The system thermalizes at higher energy density and weaker disorder, exhibiting a conducting phase. MBL violates the eigenstate thermalization hypothesis (ETH), which posits that each eigenstate at finite energy density has the same expectation values of local physical observables as the ones provided by local thermal states. This implies that thermalization occurs at the level of individual eigenstates in an isolated quantum system, where no memory of the initial state survives. The eigenstates of the system in the MBL phase do not obey the ETH, and some memory of the initial conditions is retained in local observables for arbitrarily long times.
@@
In general, the theoretical description of the MBL transition is extremely challenging. In contrast to the metal-insulator transitions, which occur in the ground state or the low-energy states of a system, the MBL transition affects the whole spectrum or at least an extended spectral range, which might be far away from the ground state. For this reason, much of the machinery to study many-body phases of matter (such as Monte Carlo techniques or tensor network methods) is not well suited for studying MBL. As a consequence, numerical explorations have mostly relied on exact diagonalization studies of systems with 20-30 spins, and even these relatively moderately sized studies strongly profit from sophisticated techniques [204, 241, 240], exploiting the sparseness of the matrix. Larger system sizes have been treated with a self-consistent Hartree-Fock description [292] or time-dependent density matrix renormalization group (DMRG) studies [153, 300].
Different fingerprints of the MBL phase are known: (i) The half-chain von Neumann entanglement entropy is strongly suppressed in the MBL phase [151, 153, 172]. The transition point can be estimated by considering the value at which the entanglement entropy per site is independent of the system size . (ii) The level spacing ratio , where is the spacing between adjacent energy levels, differentiates between localized and ergodic phase. It is well established by random matrix theory that the level spacing distribution characterizes the statistical ensemble to which a matrix belongs [105]. Ergodic systems typically belong to the Wigner-Dyson ensemble, characterized by level repulsion and a relatively large value . On the other hand, integrable/localized systems are characterized by the Poisson statistics, , see Pal and Huse [200]. (iii) Participation ratio and fractal dimension of the eigenstates, as introduced in the Section on non-interacting systems, can also be used for a many-body system if, instead of the full many-body eigenstates, the eigenstates of the one-body density matrix are considered [26, 172]. (iv) Dynamics probes have been used in many experiments: If a system is localized, it should keep memory of its initial state during a quench. Often, systems are prepared in a state with alternating empty and occupied sites, and the evolution of this even/odd population imbalance is tracked. In contrast to an ergodic system, the MBL system remains imbalanced on long time scale [232].
@@
Given the computational difficulties, quantum simulation has become an extremely valuable tool for studying MBL. At the same time, due to the instability of the MBL phase against a thermal bath, the study of the MBL phase is also motivated by the unique opportunity of isolating quantum systems from the environment offered by quantum simulators. The primary systems where the MBL transition is experimentally studied are cold atoms in optical lattices with an AA modulation [232, 169, 137, 167, 215]. Experiments in Munich [232, 169, 137] have concentrated on rather larger systems, arrays of 1D tubes with approximately 150-200 sites per tube, filled with spin-1/2 fermions.
@@
An important figure of merit in these experiments is the population imbalance between even and odd sites. Initially, the system was prepared for a large imbalance, which was then washed out through dynamic evolution. While in the absence of a quasiperiodic potential, the imbalance fully disappears on short time scales, a sufficiently strong quasiperiodic potential leads to saturation of the imbalance at a finite value, see Fig. 9a. By plotting the saturation value versus the potential strength , see Fig. 9b, the onset of localization is found to be shifted towards larger values of by the interactions. In the non-interacting case, the transition seems to occur slightly below the single-body critical value , which can be accounted for by the harmonic trapping potential compared with exact numerical simulations. In the vicinity of the localization transition, especially in the interacting system, the saturation occurs only on much longer time scales than deeply in the ergodic or the MBL regime. Kohlert et al. [137] also studied MBL in a GAA model. In contrast to the AA model, the non-interacting GAA model exhibits an intermediate phase in which both localized and delocalized eigenstates co-exist. While the experiment confirmed the MBL behavior of the interacting GAA, a many-body counterpart of the intermediate phase has not been seen.
@@
The experiments in Harvard, in particular Rispoli et al. [215], have concentrated on this critical regime, measuring observables like correlations between sites, entropy, and particle number fluctuations. The experimental procedure of Lukin et al. [167], Rispoli et al. [215] starts from a bosonic Mott insulator with one atom per site and measures the particle number per site after some evolution time. From this, different quantities can be computed. For instance, the probability of finding particles on a site determines the single-site von Neumann entropy, . For weak quasiperiodic potentials, the obtained value matches well with the expected thermal average, but it becomes reduced when the quasiperiodic modulation becomes stronger, approximately for , see Fig. 9c, notably beyond the critical value of the single-particle AA model. In contrast, the entropy reaches a maximum within the transition regime as the system strongly fluctuates. Criticality is also characterized by strong -site correlations , shown in Fig. 9(d).
@@
Bosonic systems have also been produced with up to 10 superconducting qubits, realizing Bose-Hubbard chains with an AA-like potential [219], and/or AA-modulated hopping [156]. All energy levels can be detected if prepared with only one or two bosons, revealing the fractal butterfly spectrum in the single-particle case. Using superconducting qubits, it has also been demonstrated that the level spacing of an interacting system is shifted towards smaller values by increasing the quasiperiodic modulation strength. By addressing bosonic systems, quantum simulators may finally help to solve the question of whether MBL of bosons exists [58].
@@ III.2 Topological phenomena
Topological quantum systems are characterized by a band structure with non-trivial topological properties, as in the case of topological insulators [211], or by a topological many-body ground state, as in the case of FQH systems and topological superconductors [187]. In all cases, the non-trivial topological properties are mathematically defined via non-local quantities, specifically non-zero Chern numbers, which are physically linked to robust properties, such as transport behavior or fractional quantum statistics of quasiparticles. Topological matter has been an intense research field in the past decades, mainly in the context of systems in “regular” geometries and driven by the search for robust quantum systems. Quasicrystalline or fractal structures and higher-dimensional lattices provide interesting new aspects to the field of topological matter, as discussed in the following.